i know there is a somewhat similar problem. it finds the rate of speed, but for this question, i need to find the time it took to travel a certain distance apart. i've been going at it for an hour or two, and i cant figure out how to convert it. please help.
QUESTION: Two marathon runners leave the starting gate, one running 12 mph and the other 10 mph. If they maintain the pace, how long will it take for them to be one-quarter of a mile apart?
i wasnt to sure but were you supposed to use the quater mile in feet or could leave it in a fraction of 1/4, or does it even matter?

Answer by venugopalramana(3286)

i know there is a somewhat similar problem. it finds the rate of speed, but for this question, i need to find the time it took to travel a certain distance apart. i've been going at it for an hour or two, and i cant figure out how to convert it. please help.
QUESTION: Two marathon runners leave the starting gate, one running 12 mph and the other 10 mph. If they maintain the pace, how long will it take for them to be one-quarter of a mile apart?
i wasnt to sure but were you supposed to use the quater mile in feet or could leave it in a fraction of 1/4, or does it even matter?
RELATIVE SPEED =12-10=2 MPH
DISTANCE OF SEPERATION =1/4 MILE
SO TIME =(1/4)/2=1/8 HOUR =7.5 MINUTES